A cliff is 20m above the ground. A boy rolls a ball off the edge with an unknown initial velocity, and its strikes the ground 3m horizontally away. Calculate the initial velocity of the ball 


1
Expert's answer
2021-01-24T14:24:27-0500

Let's first find the time that the ball takes to reach the ground:


"h=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2h}{g}}=\sqrt{\dfrac{2\cdot20\ m}{9.8\ \dfrac{m}{s^2}}}=2.0\ s."

Finally, we can find the initial velocity of the ball:


"x=v_0t,""v_0=\dfrac{x}{t}=\dfrac{3\ m}{2.0\ s}=1.5\ \dfrac{m}{s}."